rosemary123:

A car accelerates from rest to a velocity of 5 m/s in 4 seconds. What is its average acceleration over this period of time?

XioGonz:

average velocity is given as change in velocity / time taken. Change in velocity can be represented as v - u ie final velocity - initial velocity In this case , v = 5 m/s u = 0 t = 4 sec a = v-u / t = 5 - 0 / 4 = 5/4 = 1.25 m/s² Hence, the acceleration will be 1.25 m/s²